40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 6 22:47:41 2021, gautam, Update, ALS, Noisy ALS ALSX_excess.pngbudget.pdf
    Reply  Thu Jan 7 19:16:11 2021, gautam, Update, ALS, Noisy ALS ALS_ERR_MON.pdfAUXnoise.pdf
       Reply  Thu Jan 7 20:07:27 2021, Koji, Update, ALS, Noisy ALS 
          Reply  Thu Jan 7 21:16:22 2021, gautam, Update, ALS, Noisy ALS 
             Reply  Thu Jan 7 23:25:19 2021, Koji, Update, ALS, Noisy ALS 
          Reply  Fri Jan 8 20:01:11 2021, gautam, Update, ALS, Noisy ALS ALSX_IR_green.pdf
Message ID: 15753     Entry time: Thu Jan 7 20:07:27 2021     In reply to: 15752     Reply to this: 15754   15756
Author: Koji 
Type: Update 
Category: ALS 
Subject: Noisy ALS 

How about resurrecting the PSL table green beat for the X arm to see if the non-fiber setup shows the same level of the freq noise (e.g. the PDH locking became super noisy due to misalignment etc).

ELOG V3.1.3-